40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Feb 22 07:53:06 2022, Yehonathan, Update, WOPO, Waking up WOPO 
    Reply  Fri Feb 25 17:09:53 2022, shruti, Update, WOPO, Waking up WOPO - green beam 
       Reply  Wed Mar 2 14:38:55 2022, Yehonathan, Update, WOPO, Waking up WOPO - some more fiddling and a plan 
          Reply  Fri Mar 4 11:49:45 2022, shruti, Update, WOPO, Waking up WOPO - attempts at readout LO_shot_noise.pdf
             Reply  Thu Mar 10 17:12:01 2022, awade, Update, WOPO, Waking up WOPO - attempts at readout 
                Reply  Wed Mar 16 12:22:44 2022, Yehonathan, Update, WOPO, Waking up WOPO - attempts at readout Moku_Noise.pdfPD_Connected_no_light.pdfDiff_Channel.pdf
                   Reply  Tue Mar 22 09:19:38 2022, shruti, Update, WOPO, Waking up WOPO - back to green path before_opt.jpgaround_waveguide.png
Message ID: 2733     Entry time: Wed Mar 16 12:22:44 2022     In reply to: 2731     Reply to this: 2735
Author: Yehonathan 
Type: Update 
Category: WOPO 
Subject: Waking up WOPO - attempts at readout 

{Shruti, Yehonathan}

Yesterday, we measured a bunch of noises.

We wanted to have as reference the Moku noise, the PDs noise, and measure the shot noise of the LO again.

Attachment 1 shows the Moku noise measured by just taking data with no signal coming in. We tried both the spectrum analyzer (SA) and the oscilloscope tools, with and without averaging, and the difference between the channels.

For some reason, the SA has a worse noise figure than the oscilloscope and the difference channel doesn't give us any special common-mode rejection. Also more averaging doesn't help much because we are already taking 1.2ms of data which is way longer than 1/RBW=0.2ms we are taking here.

From now on we use the oscilloscope as the spectrum analyzer and to its noise we refer as the Moku noise floor.

Moving on, we try to measure the PD dark noise. Given that the PD dark noise floor is ~ 6nV we don't expect to see it with the Moku without amplification. Attachment 2 shows that indeed we couldn't resolve the PD dark noise.

We then opened the LO shutter. We measured with a power meter 1mW and 1.15mW coming impinging on the PDs. The voltage readings after the preamp were 1.66V for the white fiber, and 1.93 V for the red fiber. These values suggest responsivities of 0.830 and 0.834 respectively.

The PDs were measured using the Moku scope and subtracted digitally with some small gain adjustment (0.93*ch1-1.07*ch2) between the channels. The result is shown in attachment 3 together with the expected shot noise level.

1. There is not enough clearance for detecting squeezing.

2. Expected shot noise level is still too high. Does the 2kohm preamp gain go all the way above 1MHz??

Attachment 1: Moku_Noise.pdf  90 kB  Uploaded Wed Mar 16 13:32:20 2022  | Hide | Hide all
Moku_Noise.pdf
Attachment 2: PD_Connected_no_light.pdf  34 kB  Uploaded Wed Mar 16 16:54:28 2022  | Hide | Hide all
PD_Connected_no_light.pdf
Attachment 3: Diff_Channel.pdf  36 kB  Uploaded Thu Mar 17 00:07:37 2022  | Hide | Hide all
Diff_Channel.pdf
ELOG V3.1.3-